OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined: t^3-7t/t^2-64. Please help...

OpenStudy (anonymous):

Are you working with real numbers

OpenStudy (anonymous):

is this the expression: \(\frac {t^3-7}{t^2-64} \) ?? if so, factor that denominator... whatever t values make that denominator zero is where the expression itself is undefined.

OpenStudy (anonymous):

Recall that you cannot divide by 0, so simply set your denominator equal to 0 and solve when this happens For example for 1/x I set x=0 so x=0 my rational equation is undefined. Another example: 1/ t-2 set t-2=0 solve for t t=2 so you have that t=2 is undefined for your rational equation

OpenStudy (anonymous):

Can you recognize what you have in the denominator? It is a difference of squares (subtracting squares) (t)^2 - (8)^2 You have for a difference of squares a^2-b^2=(a+b)(a-b) So your a=t and b=8 so you have (t+8)(t-8) so you set this equal to 0 (t+8)(t-8)=0 What does that mean? You have two numbers that when you multiply them you get 0. So that means either one number is equal to 0 or both of them are 0. That means, t+8=0 or t-8=0 solve for t in each what do you get? t=-8 0r t=8 so when t=-8, 8 your rational expression is undefined.